OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomores tickets totaled $410. How many sophomores bought tickets? 40 90 110 120

OpenStudy (anonymous):

120

OpenStudy (anonymous):

using simultaneous eq's...let x = freshmen and...let y = sophmores therefore:----> x +y =150 and 2x +3y=410 multiply the top equation by 2 so that the x terms (ie freshmen) cancel out and the y terms (ie sophmores) are left because thats the number that you want to find... so your new set of equations will be: 2x + 2y =300 -2x + 3y =410 I subtracted the equations because i wanted to get rid of the x's (ie freshmen) therefore you get:---> -y=-110 multiplying both sides by -1 to get y...we get that y =110